(a) About how many microorganisms are found in the human intestinal tract? (A typical bacterial length scale is one micron 10- m. Assume the intestinal volume is 7 m with a diameter of 4 cm and assume bacteria occupy one hundredth of it.) 1010 O 1014 1017 O 1020 1023 0 (b) Discuss your answer to part (a). Are these bacteria beneficial, dangerous, or neutral? What functions could they serve? This answer has not been graded yet Need Help? Read It
Answer #1

(a)

Given that,

Length, L = 7 m

diameter d = 4 cm

radius, r = d / 2 = 2 cm = 0.02 m

Volume of intestinal tract, V = pi*r^2*L

V = pi*(0.02)^2 * 7 = 0.00879 m^3

Volume occupied by bacteria, V' =  0.00879 / 100 \approx 9*10-5 m63

Volume of bacteria, Vb = (10-6)3 = 10-18 m^3

Number of microorganisms,

n = V' / Vb = 9*10-5 /  10-18

n \approx 1014

(b)

These bacteria are beneficial . they live as symbiotic association with human intestine.

They help in balancing pH of intestine and digesting food. They synthesize vitamins,, amino acids and metabolize various

substances in food.